# Transfer learning from first-principles calculations to experiments with chemistry-informed domain transformation

### v1

@@ C.1 Function estimation
Figure 12 show the regression lines for each sources selected from strategy S1 with Wang2023 target data. That from the strategy S2 has been already shown in Figure 6.
@@ C.3 Target training
Figure 17 shows the target training curves for each retraining methods of the full scratch, the fine tuning, and the transfer learning, explained in Section 4.3.5. Here, C2O1 dataset is used as an example. The mean and the standard deviation of validation losses obtained by 10-fold cross validation are displayed. With the prediction models, we stopped training after 125 epochs.
@@
We also performed 10-fold cross validation during the target training with the dummy dataset. Figure 18 shows its target training curve using the C2O1 source data. To compare with the Wang2023 containing 141 target data, we limit the size of dummy data to 200 instances by random sampling. As a result, Figure 17 and 18 show similar behavior.
@@
Using full dummy data, we can investigate the behavior of the target training with a large target dataset, as shown in Figure 19. According to the result, the full-scratch and the FT exceed the TL in the case with a large dataset; It is reasonable because the last one has a lesser capacity for fixed parameters than others. However, as show in the main text, there is little difference between FT and TL in the range with O(100) data, and both are superior to the full-scratch model.

### v2

@@ C.1 Function estimation
Figure 12 shows the regression lines for each source selected from strategy S1 with Wang2023 target data. That from the strategy S2 has been already shown in Figure 6.
@@ C.3 Target training
Figure 17 shows the target training curves for each retraining method of the full scratch, the fine tuning, and the transfer learning, explained in Section 4.3.5. Here, C2O1 dataset is used as an example. The mean and the standard deviation of validation losses obtained by 10-fold cross validation are displayed. With the prediction models, we stopped training after 125 epochs.
@@
We also performed 10-fold cross-validation during the target training with the dummy dataset. Figure 18 shows its target training curve using the C2O1 source data. To compare with the Wang2023 containing 141 target data, we limit the size of dummy data to 200 instances by random sampling. As a result, Figure 17 and 18 show similar behavior.
@@
Using full dummy data, we can investigate the behavior of the target training with a large target dataset, as shown in Figure 19. According to the result, the full-scratch and the FT exceed the TL in the case with a large dataset; It is reasonable because the last one has a lesser capacity for fixed parameters than the others. However, as shown in the main text, there is little difference between FT and TL in the range with O(100) data, and both are superior to the full-scratch model.